# Linear-time classical approximate optimization of cubic-lattice classical spin glasses

### v4

@@ Abstract
Computing low-energy configurations (i.e., approximate optimization) of classical spin glasses is of relevance to both condensed matter and combinatorial optimization. Recent theoretical work opens the possibility to make its time complexity with quantum annealing generically polynomial, and D-Wave experiments can now achieve approximate optimization of cubic-lattice classical Ising spin glasses with spins. It is therefore timely to ask which short-range classical spin glasses are good candidates for demonstrating quantum advantage in the time complexity of heuristic approximate optimization. One intuition is to consider models with very rugged energy landscapes in configuration space. However, here we provide evidence that short-range classical spin glasses may be approximately optimized in linear time and space with a very simple deterministic tensor-network heuristic regardless of ruggedness. On the cubic lattice with up to 505050 spins, we obtain energy errors of 3% for the model used in recent D-Wave experiments, and 5% for much more rugged planted-solution instances. For cubic-lattice-Ising reductions of unweighted Max-Cut on random 3-regular graphs with up to 300 vertices, we find energy errors of 1% and approximation ratios of about 72-88%. A theoretical implication of our algorithm is that only quantum algorithms with sublinear time complexity may have time-complexity quantum advantage for heuristic approximate optimization of classical spin glasses that have decaying correlations. These results inform the search for quantum advantage and suggest an efficient classical method for generating warm starts for other spin-glass optimization algorithms. Our algorithm is amenable to massive parallelization and may also allow for low-power, accelerated implementations with photonic matrix-multiplication hardware.
@@ I Introduction
Classical, short-range spin-glass Hamiltonians have both fundamental and applied significance. One example is the classical Ising model on the simple cubic lattice, whose Hamiltonian is given by
@@
where , , and are lattice-site indices, and denotes a restriction to nearest neighbors on the simple cubic lattice. When follows a random distribution (typically bimodal or Gaussian), this model is known as the classical Edwards-Anderson Ising model with zero field and is thought to capture the fundamental physics of spin-glass materials [1]. Yet, its low-temperature phase remains incompletely understood despite decades of effort [2]. On the applied side, computing the ground state of Eq. (1) is known to be NP-hard [3], and a more structured choice of allows it to encode problems in discrete combinatorial optimization [4]. Therefore, computing the low-energy solutions (i.e., approximate optimization) of Eq. (1) can be useful for both condensed matter physics and combinatorial optimization.
Recent developments in quantum annealing make the simple-cubic-lattice model of Eq. (1) particularly relevant to the issue of quantum advantage: On the experimental front, D-Wave and collaborators have recently demonstrated the approximate optimization of such a model with over 5,000 spins with pure quantum annealing [5] and over 11,000 spins with hybrid quantum annealing [6]. On the theoretical side, the recent work in Refs. [7, 8] presents the possibility of achieving generic polynomial time complexity for heuristic approximate optimization 111by “heuristic approximate optimization” we mean an approximate minimization of the cost function (i.e., Hamiltonian) in a way that lacks a formal guarantee on the error of the minimization of classical Ising spin glasses with quantum annealing. Therefore, quantum advantage in the computational time complexity for heuristic approximate optimization of Eq. (1) may be possible in the near- to mid-term, but it is not clear for which instance classes of , if any; understanding where known classical algorithms are slow with this task can provide clues.
@@
State-of-the-art Markov-Chain-Monte-Carlo algorithms are considered to be some of the best general classical heuristics for approximate and exact optimization of spin glass models. Consequently, (classical) computational hardness has become synonymous in the spin-glass literature with Monte-Carlo hardness (e.g., Ref. [10]). For this reason, Ref. [11] suggests that spin-glass models that display a steep increase in the autocorrelation time of parallel-tempering Monte Carlo as the temperature is lowered are good places to demonstrate quantum advantage. Also, Ref. [12] shows a time-complexity advantage of (error-corrected) quantum annealing over parallel tempering Monte Carlo (with isoenergetic cluster moves) for approximate optimization of a quasi-2D spin glass, and concludes this to be a demonstration of algorithmic quantum speedup. The hardness of certain spin glass instances for Monte-Carlo (and other) optimization algorithms that traverse the spin-configuration space is understood to be due to the rugged energy landscape (i.e., a plethora of local minima) that arises at low temperatures in that space. Even machine-learning enhancements of Monte Carlo methods have been shown to be ineffective at overcoming this source of computational hardness [13]. However, it is thought that quantum algorithms may be able to overcome this source of hardness via quantum tunneling [14, 15, 16].
It would seem, then, that one simply needs to find instances of Eq. (1) whose approximate optimization with the best Monte-Carlo algorithms is computationally hard (i.e., exhibits or at least hints at superpolynomial time complexity) in order to identify good candidates for demonstrating time-complexity quantum advantage in heuristic approximate optimization. However, Ref. [17] provides evidence that short-range rugged-energy-landscape spin glasses may be heuristically approximately optimized classically in quadratic time with approximate contractions of tensor networks, though the wall time for cubic lattices with that heuristic proves to be too large to be practical for large systems. Here we present a major advance over that result: We demonstrate a very simple deterministic tensor-network heuristic that is able to compute in linear time (and space) at least some (but not necessarily all) spin configurations that have energies within a few percent of the ground-state energy for the following three versions of Eq. (1): (1) the cubic-lattice model that is used in the D-Wave approximate optimization experiments in Refs. [5] and [6], (2) a planted-solution, cubic-lattice model whose hardness (i.e., time to exact optimization) for state-of-the-art Monte Carlo heuristics is orders of magnitude more than the model, and (3) a cubic-lattice model that encodes (unweighted) Max-Cut on random 3-regular graphs. Unlike the heuristic in Ref. [17], the present one is able to address very large cubic-lattice systems in a practical amount of time. We also provide an explanation for why this heuristic implies that a quantum algorithm must have sublinear time complexity in order to have quantum advantage in the time complexity of heuristic approximate optimization of classical spin glasses that have correlations that decay with euclidean distance, such as models (1) and (2) just mentioned.
@@
The tensor-network algorithm presented here relies on a tensor-network representation of the partition function, and is made possible through the following line of developments: First, the work in Refs. [18, 19, 20, 21] presented ways of using tensor-network algorithms for homogeneous classical lattice models. This included a method of sampling from the Boltzmann distribution of homogeneous, two-dimensional classical spin lattices [21]. Then, Ref. [22] showed that the partition function of an inhomogeneous classical spin lattice may be exactly represented by the contraction of a network of tensors where the geometry of the network reflects that of the Hamiltonian’s interaction graph; while an exact contraction of the full network has an exponential cost, the contraction of the full network may be approximated, via truncated matrix decompositions, in polynomial time. Ref. [23] then, analogously to the work in Ref. [21] for homogenous systems, demonstrated how to use this idea to sample (via computation of conditional marginals) the low-temperature Boltzmann distribution of planar and quasi-planar spin glasses in polynomial time. Finally, Ref. [17] used a technically different type of approximate contraction from Ref. [23] to generate near-optimal solutions of spin glasses on periodic square and cubic lattices in quadratic time. In both of the latter two works, the time complexity is determined by the complexity of contracting the network and not the Monte-Carlo hardness (which is related to the ruggedness of the free-energy landscape in configuration space), and the heuristic is successful in generating near-optimal solutions for all tested problem instances. We note, however, that the data in Ref. [17] for cubic-lattice spin glasses only spans two sizes (444 and 666) due to the high absolute time cost.
@@
Compared to the approach in Ref. [17], here we present a more efficient way of using tensor-network representations of classical, cubic-lattice spin-glass partition functions for approximate optimization. In Refs. [23] and [17], the amount of the network that is utilized to compute conditional marginals for single spins grows with the number of spins () that are conditioned upon (though Ref. [17] always uses the full network for simplicity). This leads to a time complexity for computing a total spin configuration that is quadratic in the total number of spins (an exponential complexity is avoided by making approximations via truncated matrix decompositions). The method in the present work instead has a linear time (and memory) complexity for short-range classical spin glasses, and also eliminates the need for matrix decompositions. These improvements result in a dramatic speedup, and can be understood via a very simple intuition about short-range-correlated spin glasses: bootstrapping of approximate local energy minimization should lead to approximate global energy minimization. By “short-range-correlated” we mean that the couplings of the short-range Hamiltonian do not have long-range correlations, and by “bootstrapping” we mean requiring consistency between the spin configurations of the overlapping system fragments that are used for the local energy minimizations. Since the approximate local energy minimizations are done via exact contractions of tensor-network fragments of fixed size, bootstrapping them over the system requires no matrix decompositions and has () time complexity. While such a local optimization method will be biased against spin configurations with large-scale structures, such as droplets, it is intuitive that not all low-energy configurations will have such structures. We find that this method allows us to approximately optimize short-range-correlated cubic-lattice classical spin glasses that are orders of magnitude larger than those in Ref. [17]. Somewhat surprisingly, we also find the method to work well for cubic-lattice Hamiltonians that encode the long-range-correlated problem of unweighted Max-Cut on random 3-regular graphs.
The present heuristic shares the local-optimization spirit of the algorithm in Sec. III. of Ref. [24], but it is different in the following crucial ways: (1) the present algorithm aims at only approximate optimization instead of exact optimization, (2) the present algorithm has a guaranteed linear time complexity, and (3) the present algorithm consists almost purely of matrix multiplications (i.e., exact contractions of tensor networks).
@@
Due to the ubiquity of matrix multiplication in computing, specialized photonic hardware for it, which seeks to improve both power consumption and speed over conventional processors, is under active development [25, 26, 27, 28]. The present heuristic thereby serves, in principle, as the basis for a new type of specialized computing machine for obtaining near-optimal solutions of certain types of discrete combinatorial optimization problems.
In Section II we provide details of the heuristic algorithm. In Sections III.1, III.2, and III.3 we respectively present results for the model; the hardest instance class of the cubic-lattice tile-planted-solution model; and the cubic-lattice model that encodes (unweighted) Max-Cut on random 3-regular graphs. In Section IV we explain the implications of the algorithm for where quantum advantage may theoretically be found. We conclude in Section V with a summary and discussion.
II algorithm
@@
As explained in Refs. [23, 17, 29] both conditional and unconditional single-spin marginals for a classical spin glass in the canonical ensemble can be computed via the contraction of a tensor network wherein the indices of a given tensor correspond to single-spin configurations and the elements of the tensor correspond to the Boltzmann weights of the joint configurations of the spins at the tensor’s legs (e.g., a two-index tensor contains the Boltzmann weights for the configurations of two spins); the geometry of the tensor network mirrors the geometry of the Hamiltonian’s interaction graph. The contraction of the network results in a multiplication of the local Boltzmann weights across the whole system that yields the (conditional or unconditional, depending on details of the network) marginal for the spin of interest. See Fig. 1 and Appendix A of Ref. [17] for further explanation. Computation of conditional single-spin marginals, and single-spin decimations according to those single-spin marginals, allows one to sample the Boltzmann distribution
@@
The algorithm in this work is a modification of the one in Ref. [17]: the marginal of a single spin is (approximately) computed by exactly contracting only a fragment of the network that is local to the spin of interest. Sequentially computing single-spin marginals with tensor-network fragments that overlap previously decimated spins results in a bootstrapping of approximate local energy minimization if is sufficiently large (see Fig. 2 for an illustration on the square lattice). The intuition behind this approach is that if the Hamiltonian is not long-range correlated, then such a bootstrapping should also result in an approximate global energy minimization. Such an outcome would be consistent with the finding in Ref. [17] that approximate optimization can be successfully accomplished for short-range-correlated Hamiltonians with approximate contractions of the full tensor network that use only small bond dimensions. In this work we implement the new algorithm with the Python libraries quimb [30] and cotengra [31] on an Apple M2 Ultra CPU (16 performance and 8 efficiency cores) with 128 GB of RAM. We note that in two and higher dimensions this algorithm is straightforward to parallelize by using multiple fragments simultaneously, though we do not do so here.
III Results

### v6

# Linear-time classical approximate optimization of rugged-energy-landscape cubic-lattice classical spin glasses

@@ Abstract
Rugged energy landscapes present computational difficulty for optimization. A quantum alleviation may be possible via tunneling. Classically, an alleviation may arise through optimizing over subsystems and concatenating the results, but this introduces error in the optimization of the full problem and thereby raises the question of whether such an approach actually has a scaling advantage over approximately optimizing the full system directly to the same error. Here we investigate this question in the setting of cubic-lattice classical Ising spin glasses, where recent theoretical and experimental developments open the possibility of showing quantum speedup with quantum annealing, and where classical time-complexity results remain absent. For the subsystem-based approach we introduce a very simple deterministic tensor-network heuristic that features linear time and space complexity for approximate optimization of the full problem. For the full-system approach we use simulated annealing and parallel tempering. For the most rugged instances generated with tile planting on system sizes up to 565656, we find that full-system simulated annealing and full-system parallel tempering display a slightly superlinear time complexity when targeting the same energy error achieved by the subsystem-based linear-time heuristic. We also find that the error of the latter heuristic monotonically decreases with increasing ruggedness over the higher end of the ruggedness spectrum. These results suggest that subsystem-based classical optimization heuristics should be taken into account when seeking to demonstrate quantum speedup on rugged energy landscapes. We discuss prospects for reducing the error of our heuristic, for adapting our heuristic to arbitrary graphs, and for low-power, accelerated implementations with photonic matrix-multiplication hardware.
@@ I Introduction
Classical spin-glass models have both fundamental and applied significance, and the simplest possible class of such Hamiltonians is the Ising spin glass in zero field, given by
@@
where , , and and are site indices. Computing the ground state of Eq. (1) is known to be NP-hard [1], and a more structured choice of allows it to encode problems in discrete combinatorial optimization [2] such that the Hamiltonian encodes a cost function to be minimized and the ground state corresponds to the optimal solution of the optimization problem. On a -dimensional lattice with only nearest-neighbor interactions, when follows a random distribution (typically bimodal or Gaussian) this model is known as the classical Edwards-Anderson Ising model with zero field and is thought to capture the fundamental physics of spin-glass materials [3]. Yet, its low-temperature phase remains incompletely understood despite decades of effort [4].
I.1 Quantum speedup for approximate optimization
While computing the ground states (i.e., exact optimization) of classical spin glasses is therefore of both fundamental and applied interest, computing their low-energy states (i.e., approximate optimization) can also be useful. The low-energy excited states of classical spin-glass Hamiltonians, for instance, can provide additional insights into the low-temperature phases of classical spin-glass materials. For combinatorial optimization, merely reduced-cost solutions can be of practical benefit, and what is often desired is a set of low-energy states that satisfy various constraints [5, 6]. Even in cases where the optimal solution is the goal, the outputs of approximate optimization algorithms could potentially serve as warm starts (i.e., initial conditions that result in a shorter runtime compared to random initial conditions) for exact optimization algorithms.
There is a long history of investigating the possibility of quantum advantage (i.e., some type of performance advantage of quantum algorithms over classical algorithms) for the optimization of classical spin glasses [7, 8]. A key performance metric is the time-to-solution (TTS). In this work we are concerned specifically with the type of quantum advantage that is termed quantum speedup [9], which means a superior (i.e., smaller) time complexity (i.e., scaling of the TTS as a function of system size ()) compared to the best-known time complexity of classical algorithms for the same computational task. An advantage in time complexity is sometimes referred to as a scaling advantage.
For exact optimization there have been a number of results showing the potential for quantum speedup [10, 11, 12, 13, 14], though an actual demonstration remains absent. Regarding quantum speedups in approximate optimization, Refs. [15, 16] theoretically show the potential for it; Ref. [17] claims an actual demonstration of quantum speedup of D-Wave quantum annealing by comparing its time complexity to that of parallel tempering with isoenergetic cluster moves (a type of Markov-Chain-Monte-Carlo algorithm) for the approximate optimization of a class of quasi-2D spin glasses, but the more recent work of Ref. [18] refutes this claim by showing that the Simulated Bifurcation Machine [19, 20], a classical heuristic inspired by nonlinear Hamiltonian dynamics, achieves better time complexity than quantum annealing in the same class of problems. Thus it is apparent that claims of quantum speedup for classical spin glass optimization should not be based on comparisons with Markov-Chain-Monte-Carlo algorithms alone.
In this work we are concerned specifically with the issue of quantum speedup for the computational task of approximate optimization of Eq. (1) on the simple cubic lattice. To the best of our knowledge, this issue has not yet been addressed in the literature: the time complexity of neither classical nor quantum algorithms has been established for this computational task. Yet, the state-of-the-art of D-Wave quantum annealing encompasses approximate optimization of such models: D-Wave has demonstrated the approximate optimization of a cubic-lattice classical Ising spin glass with over 5,000 spins with pure quantum annealing [21] and over 11,000 spins with hybrid quantum annealing [22]. Experimental investigation of the time complexity of quantum annealing for this computational task is therefore feasible. Further, recent theoretical results [23, 24, 25, 26, 27] suggest the possibility of modifying quantum annealing so as to achieve a substantial improvement in the time complexity for this computational task compared to ordinary quantum annealing. It is therefore timely to ask for data on the time complexity of classical heuristics for the approximate optimization of cubic-lattice versions of Eq. (1); here we address this by presenting and testing a new classical heuristic that is tailored to such models, and also by providing a comparison of that new heuristic with the Markov-Chain-Monte-Carlo algorithms simulated annealing (SA) and parallel tempering (PT). Our results reveal that the new heuristic at least sometimes has a scaling advantage over SA and PT for approximate optimization of cubic-lattice classical Ising spin glasses. Thus, any attempts to show quantum speedup on this class of problems must take this new heuristic, and potential improvements thereof, into account.
I.2 Energy-landscape ruggedness as a source of computational difficulty
@@
State-of-the-art Markov-Chain-Monte-Carlo (MCMC) algorithms, such as SA and PT, are often considered to be the best general classical heuristics for approximate and exact optimization of classical spin-glass models [17] (but see Ref. [18]). Consequently, computational hardness has become synonymous in the spin-glass literature with MCMC hardness (e.g., Ref. [28]). For this reason, Ref. [29] suggests that spin-glass models that display a steep increase in the autocorrelation time of parallel-tempering Monte Carlo as the temperature is lowered are good places to look for quantum advantage in optimization. The optimization hardness of certain spin glass instances for MCMC algorithms is understood to be due to the rugged energy landscape (i.e., a plethora of local minima) in spin-configuration space [30, 31]. Moreover, machine-learning enhancements of Monte-Carlo methods have been shown to be ineffective at overcoming this source of computational hardness [32], and non-MCMC classical heuristics that model dynamical systems evolution, such as the Simulated Bifurcation Machine mentioned previously, also display a substantial performance drop with increasing landscape ruggedness [33, 34]. In contrast, it is thought that quantum algorithms (such as quantum annealing) may be able to (at least partially) overcome this source of hardness via quantum tunneling [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45].
In this work we make the simple observation that a potential classical way of partially overcoming the difficulty of approximately optimizing over rugged energy landscapes is to optimize over subsystems and simply concatenate the results to approximate the solution to the full problem; this is a linear time heuristic. The intuition here is that if much or all of the energy landscape is very rugged, then efficiently arriving at a low-energy solution in this manner will bypass much of the ruggedness. Indeed, our results in this work do show a scaling advantage for our subsystem optimization-based heuristic compared to full-system MCMC on extremely rugged energy landscapes when targeting a solution that is within a few percent relative energy error of the ground state. While approximately solving larger optimization problems via optimizing their subsystems is not a new idea, to our knowledge this idea has not yet been investigated specifically as a means of alleviating the computational difficulty presented by energy-landscape ruggedness.
I.3 A subsystem optimization-based heuristic for cubic-lattice spin glasses
@@
Our subsystem-based classical heuristic relies on a tensor-network representation of the partition function, and is made possible through the following line of developments: First, the work in Refs. [46, 47, 48, 49] presented ways of using tensor-network algorithms for homogeneous classical lattice models. This included a method of sampling from the Boltzmann distribution of homogeneous, two-dimensional classical spin lattices [49]. Then, Ref. [50] showed that the partition function of an inhomogeneous classical spin lattice may be exactly represented by the contraction of a network of tensors where the geometry of the network reflects that of the Hamiltonian’s interaction graph; while an exact contraction of the full network has an exponential cost, the contraction of the full network may be approximated, via truncated matrix decompositions, in polynomial time. Ref. [51] then, analogously to the work in Ref. [49] for homogeneous systems, demonstrated how to use this idea to sample (via computation of conditional marginals) the low-temperature Boltzmann distribution of planar and quasi-planar spin glasses in polynomial time. Finally, Ref. [52] used a technically different type of approximate contraction from Ref. [51] to generate near-optimal solutions of spin glasses on periodic square and cubic lattices in quadratic time. In both of the latter two works, the time complexity is determined by the complexity of contracting the network and not the Monte-Carlo hardness (which is related to the ruggedness of the free-energy landscape in configuration space), and the heuristic is successful in generating near-optimal solutions for all tested problem instances. We note, however, that the data in Ref. [52] for cubic-lattice spin glasses only spans two sizes (444 and 666) due to the high absolute time cost.
@@
Compared to the approach in Ref. [52], the new method we present here constitutes a more efficient way of using tensor-network representations of classical, cubic-lattice spin-glass partition functions for approximate optimization. In Refs. [51] and [52], the amount of the network that is utilized to compute conditional marginals for single spins grows with the number of spins () that are conditioned upon (though Ref. [52] always uses the full network for simplicity). This leads to a time complexity for computing a total spin configuration that is quadratic in the total number of spins (an exponential complexity is avoided by utilizing truncated matrix decompositions). The method in the present work instead has a linear time (and memory) complexity for short-range classical spin glasses, and also eliminates the need for matrix decompositions. These improvements result in a dramatic reduction in walltime, and can be understood via a very simple intuition about short-range-correlated spin glasses: bootstrapping of approximate local energy minimization should lead to approximate global energy minimization. By “short-range-correlated" we mean that the short-range Hamiltonian does not amount to an embedding of a long-range Hamiltonian (as in, for example, Ref. [53]). Therefore, in contrast to the method in Ref. [52], the one presented here is able to reach large system sizes on the cubic lattice and thereby provide substantial data to challenge the time complexity of quantum algorithms for the approximate optimization of rugged-energy-landscape classical spin glasses on the cubic lattice.
@@
Also in contrast to the tensor-network method in Ref. [52], the absence of matrix decompositions from the new method in this work means that its computational cost is dominated by matrix multiplications and it can therefore be substantially sped up with GPUs or FPGAs. Further, due to the ubiquity of matrix multiplication in computing, specialized photonic hardware for it, which seeks to improve both power consumption and speed over conventional processors, is under active development [54, 55, 56, 57]. The new heuristic thereby serves, in principle, as the basis for a new type of specialized computing machine for obtaining near-optimal solutions of certain types of discrete combinatorial optimization problems.
The approximate local energy minimizations in the new method are done via exact contractions of tensor-network fragments of fixed size, which is why bootstrapping them over the system requires no matrix decompositions and has () time complexity. While such a local optimization method will be biased against spin configurations with large-scale structures, such as droplets, it is intuitive that not all low-energy configurations will have such structures. However, small perturbations to the bonds in classical cubic-lattice spin glasses can substantially alter the ground state [58, 59], so it is not a priori obvious how small of an average error such a bootstrapping heuristic will yield, nor is it obvious how much fluctuation will be in that error from instance to instance within a fixed hardness class. Our results from all of the levels of energy-landscape ruggedness that we test reveal an upper bound on the average relative energy error of and a variance that varies widely over the different instance classes. The average-error upper bound is substantially smaller than the smallest upper bound of that is theoretically proven (assuming PNP) for the error of non-heuristic polynomial-time approximation algorithms for Eq. (1) (see End Matter of Ref. [60]).
The new heuristic presented here shares the local-optimization spirit of the algorithm in Sec. III. of Ref. [61], but it is different in the following crucial ways: (1) the present algorithm aims at only approximate optimization instead of exact optimization, (2) the present algorithm has a guaranteed linear time complexity, and (3) the present algorithm consists almost purely of matrix multiplications (i.e., exact contractions of tensor networks).
I.4 Outline for the rest of the work
In Section II we provide details of our tensor-network heuristic. In Section III.1 we present results from our heuristic for the cubic-lattice model (the model that is used in recent D-Wave experiments [21, 22]). In Section III.2 we present results from our heuristic on two classes of the cubic-lattice tile-planted-solution model across their full range of energy landscape ruggedness, and we present results from simulated annealing and parallel tempering on the most rugged instances of that same model. We conclude in Section IV with a summary and outlook.
II Tensor-Network heuristic details
@@
As explained in Refs. [51, 52, 62] both conditional and unconditional single-spin marginals for a classical spin glass in the canonical ensemble can be computed via the contraction of a tensor network wherein the indices of a given tensor correspond to single-spin configurations and the elements of the tensor correspond to the Boltzmann weights of the joint configurations of the spins at the tensor’s legs (e.g., a two-index tensor contains the Boltzmann weights for all the possible configurations of two spins); the geometry of the tensor network mirrors the geometry of the Hamiltonian’s interaction graph. The contraction of the network results in a multiplication of the local Boltzmann weights across the whole system that yields the (conditional or unconditional, depending on details of the network) marginal for the spin of interest. See Fig. 1 and Appendix A of Ref. [52] for further explanation. Computation of conditional single-spin marginals, and single-spin decimations according to those single-spin marginals, allows one to sample the Boltzmann distribution
@@
The algorithm in this work is a modification of the one in Ref. [52]: the marginal of a single spin is (approximately) computed by exactly contracting only a fragment of the network that is local to the spin of interest. Sequentially computing single-spin marginals with tensor-network fragments that overlap previously decimated spins results in a bootstrapping of approximate local energy minimization if is sufficiently large (see Fig. 2 for an illustration on the square lattice). The intuition behind this approach is that if the Hamiltonian is not long-range correlated, then such a bootstrapping should also result in an approximate global energy minimization. Such an outcome would be consistent with the finding in Ref. [52] that approximate optimization can be successfully accomplished for short-range-correlated Hamiltonians with approximate contractions of the full tensor network that use only small bond dimensions. In this work we implement the new algorithm with the Python libraries quimb [63] and cotengra [64] on an Apple M2 Ultra CPU with 16 performance cores, 8 efficiency cores, and 128 GB of RAM. We note that in two and higher dimensions this algorithm is straightforward to parallelize by using multiple fragments simultaneously, though we do not do so here. Appendix A contains details of the SA and PT implementations.
III Experiments